OpenStudy (anonymous):

In order to solve the equation -3x - 6 = 5, the first step is to: add 6 to both sides divide by 5 on both sides subtract 6 from both sides multiply by 5 on both sides

OpenStudy (fibonaccichick666):

what do you think?

OpenStudy (anonymous):

I add 6 to both sides.

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

-2x - 15 = 6x + 9 3 -3 -1 6

OpenStudy (fibonaccichick666):

what do you think?

OpenStudy (anonymous):

Idkk im confused.

OpenStudy (fibonaccichick666):

so what is your first step?

OpenStudy (anonymous):

i think you add 15 to both sides but idk.

OpenStudy (fibonaccichick666):

that works

OpenStudy (fibonaccichick666):

go ahead, and what do you get?

OpenStudy (anonymous):

do you add it to 15 and 6?

OpenStudy (fibonaccichick666):

? I don't understand that comment

OpenStudy (anonymous):

Like what are you adding 15 to?

OpenStudy (anonymous):

the 15 and 6 right?

OpenStudy (fibonaccichick666):

both sides, you don't add it to anything in particu;lar, just tack it on the end

OpenStudy (fibonaccichick666):

ie 4v+7=12 4v-7+7=12-7

OpenStudy (anonymous):

woah where did that come from lol so far i think i have it but ill show you -2x-30=21x+9 -32x=21x+9 -32=30x?

OpenStudy (anonymous):

Idkim so confused.

OpenStudy (fibonaccichick666):

not quite

OpenStudy (anonymous):

ofcourse not lol.

OpenStudy (fibonaccichick666):

-2x - 15 = 6x + 9 so add 15 is good, but you do it like this \[-2x - 15 = 6x + 9\]\[-2x-15+15=6x+9+15\]

OpenStudy (fibonaccichick666):

now simplify

OpenStudy (anonymous):

so the first half is -2?

OpenStudy (fibonaccichick666):

no, you can't just drop the variable,

OpenStudy (fibonaccichick666):

make sure you keep it

OpenStudy (anonymous):

omgg that is dumb. So -2x=30x? is that wrong. or what im about to give up

OpenStudy (fibonaccichick666):

you have the left, now lets work on the right. so we need to simplify(aka combine like terms) 6x+9+15

OpenStudy (fibonaccichick666):

so let me make it a little more obvious: \[6x+(9+15)\]

OpenStudy (fibonaccichick666):

^I can do that because addition is associative ^

OpenStudy (anonymous):

so then itll be 6x+24?

OpenStudy (fibonaccichick666):

yep, so now you have \[-2x=6x+24\]

OpenStudy (anonymous):

yea

OpenStudy (fibonaccichick666):

what do you think you do now?

OpenStudy (anonymous):

subtract -2?

OpenStudy (fibonaccichick666):

no, remember we are getting all of the like terms together, so how could you get the 6x to the left side?

OpenStudy (anonymous):

umm, maybe adding the like terms?

OpenStudy (anonymous):

or subtracting

OpenStudy (fibonaccichick666):

well, if I added 6x to both sides, what would I get? how about if I subtract?

OpenStudy (anonymous):

-2x+6x=4 -2x-6x=-8

OpenStudy (anonymous):

oh wait haha read that wrong. i think.

OpenStudy (fibonaccichick666):

what happened to the 24 on the right?

OpenStudy (anonymous):

-2x+24 = 22 =2x-24=-26

OpenStudy (anonymous):

right?

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

stahp

OpenStudy (fibonaccichick666):

ok, you have the left correct two steps back, but you changed the right. Don't it should be 24

OpenStudy (anonymous):

would x = 18?

OpenStudy (fibonaccichick666):

and you cannot add something with an x in it and something without an x

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

go back to here

OpenStudy (fibonaccichick666):

-2x+6x=4 -2x-6x=-8

OpenStudy (anonymous):

omgggg i hate math.

OpenStudy (fibonaccichick666):

now, change the right to 24

OpenStudy (fibonaccichick666):

you just have not learned the rules yet

OpenStudy (anonymous):

what do you mean change it to 24?

OpenStudy (fibonaccichick666):

I mean this: -2x+6x=12x+24 -2x-6x=24

OpenStudy (fibonaccichick666):

now, which is simpler?

OpenStudy (anonymous):

so now wouldnt you subtract the -2x-6x

OpenStudy (fibonaccichick666):

but you retain the x when you do

OpenStudy (anonymous):

yes so it would be 4x=24?

OpenStudy (fibonaccichick666):

no, what is (-2-6)x=24

OpenStudy (fibonaccichick666):

simplify the parentheses

OpenStudy (anonymous):

-8

OpenStudy (fibonaccichick666):

so then we get -8x=24

OpenStudy (anonymous):

now what.

OpenStudy (anonymous):

divide each side by -8?

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

x=-3?

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

omggg

OpenStudy (fibonaccichick666):

please do yourself a favor... Khanacademy.org go there

OpenStudy (fibonaccichick666):

medal please?